OpenStudy (anonymous):

What is the solution to the system of equations? -2x+2y-2=10 -3x+y-z=5 2x-3y+z=-1

OpenStudy (jack1):

@Loser66 i made the assumption it was a typo in 2nd post

OpenStudy (anonymous):

no, my answers are either: (1,-2,-7) (0,-2,7) (0,-2,-7) (0,2,7)

OpenStudy (jack1):

check ur first equations in the original question, is what you typed correct? or should there be a z in the first line?

OpenStudy (anonymous):

z in the first line

OpenStudy (jack1):

so it should be: -2x+2y- 2Z =10 -3x+y-z =5 2x-3y+z =-1 ?? yeah?

OpenStudy (anonymous):

yep that is correct

OpenStudy (loser66):

0,-2,-7 seems good, to me, but don't trust me, wait for Jack1

OpenStudy (loser66):

@Jack1 I'm cheating on the answers, hihhi

OpenStudy (jack1):

-2x+2y- 2Z =10 -3x+y-z =5 2x-3y+z =-1 do you understand how to solve simultaneous equations? so using eqn 3, we can rearrange to form : z =-1 + 3y - 2x then we can put that into eqn 1 -2x+2y- 2(-1 + 3y - 2x) =10 so -2x +4x +2y - 3y +2 = 10 so 2x -y = 8 so y = 2x - 8 sub y = 2x - 8 into eqn 2 with z =-1 + 3y - 2x and you'll get an answer for x @Loser66 is correct, you should end up with z = -7, y = -2, x = 0

OpenStudy (loser66):

@Jack1 when we don't have options to choose, we have to solve. When we have them , cheat on it, why not?

OpenStudy (jack1):

@Loser66 Give a fish a fish, and you introduce the concept of slavery to an entire species. Teach a fish to fish, and you make him a mythical, nightmarish murderer feared by his own kind. So maybe just stay away from fish, ok?

OpenStudy (jack1):

;)

OpenStudy (loser66):

yes, my bad, I apology

OpenStudy (anonymous):

thank you Jack1 and Loser66